Find exact value of sin^-1 (root3/2)
Answer by Alan3354(69443)

Find exact value of sin^-1 (root3/2)
---------------
= 60 degs, 120 degs + n*360 degs, n = 0,1,2,3...
or = pi/3, 2pi/3 + n*2pi, n = 0,1,2,3...